# Model-Based Diffusion Sampling for Predictive Control in Offline Decision Making

### v1

@@ 4.2 Constrained Offline Decision Making
Results on Standardized Benchmarks (DSRL): We evaluate on the DSRL benchmark (dsrl), which includes safety-critical velocity and Safety Gym tasks. The objective is to maximize return while keeping cumulative cost below a specified budget. We compare against behavior cloning (BC-All), behavior cloning trained only on safe trajectories (BC-Safe), cost-regularized approaches such as COptiDICE (lee2022coptidice), as well as transformer-based CDT (liu2023constrained). Our method, MPDiffuser, is tested using 16 samples with learned cost and reward functions parameterized as MLPs. For each task, we evaluate all methods under cost budgets of $20$, $40$, and $80$, reporting average normalized return and cost over $60$ independent trials per budget.
@@
BC-All COptiDICE BC-Safe CDT MPDiffuser
Return Cost Return Cost Return Cost Return Cost Return Cost
HopperVelocity 0.65 6.39 0.13 1.51 0.36 0.67 0.63 0.61 0.81 0.37
Walker2dVelocity 0.79 3.88 $0.12$ $0.74$ $0.79$ $0.04$ $0.78$ $0.06$ 0.80 0.27
HalfCheetahVelocity 0.97 13.1 $0.65$ $0.0$ $0.88$ $0.54$ 1.0 0.01 0.98 0.77
PointGoal1 $0.65$ $0.95$ 0.49 1.66 $0.43$ $0.54$ 0.69 1.12 0.74 0.88
PointCircle1 0.79 3.98 0.86 5.51 $0.41$ $0.16$ 0.59 0.69 0.58 0.94
CarGoal1 $0.39$ $0.33$ $0.35$ $0.54$ $0.24$ $0.28$ 0.66 1.21 0.63 0.92
CarCircle1 0.72 4.39 0.7 5.72 0.37 1.38 0.6 1.73 0.50 0.85
Table 3: Performance on DSRL benchmark tasks. Return–cost tradeoffs on the DSRL benchmark (dsrl). Unconstrained baselines (BC-All) achieve high rewards but violate safety constraints with excessive costs, while cost-regularized methods (COptiDICE, BC-Safe) sacrifice performance for safety. MPDiffuser achieves competitive returns while maintaining safety, demonstrating effective safety-performance balance.
Importance of Dynamic Feasibility for Ranker: We evaluate our approach on the classic Pendulum environment, modified with a hard velocity constraint requiring angular velocity to remain below $6.5\,\mathrm{m/s}$. We first train a standard soft actor-critic (SAC) (haarnoja2018soft) agent for 100100k steps, which frequently violates the velocity constraint, and a safe SAC agent that penalizes constraint violations heavily. To construct the dataset, we use the replay buffer of the unsafe SAC agent and $300$ trajectories collected from the safe SAC agent. We then compare MPDiffuser with only planner based sampling and SafeDiffuser (xiao2023safediffuser) enforces safety by projecting sampled states within constraints during the diffusion process using a barrier-function-based approach.
Table 4 shows success rates as a function of sampled trajectories, where success means stabilizing the pendulum upright without violating the velocity constraint. SafeDiffuser achieves substantially lower success rates than other methods, underscoring the necessity of dynamic feasibility for projection-based approaches to maintain safety. While planner initially improves with more samples, its performance later drops significantly because it often generates dynamically infeasible rollouts. With more samples, the chance of selecting such “hallucinated” trajectories that appear high-return but fail in practice increases. In contrast, MPDiffuser sustains high success rates as samples grow, highlighting robustness induced by our alternating sampling scheme.
@@ 4.3 Extending MPDiffuser to Visual Domains
Figure 8: Illustration for visual inputs.
@@
To assess whether our framework can be extended to high-dimensional visual inputs, we conduct a preliminary proof-of-concept experiment. Specifically, we consider the Pendulum environment with image observations. We first train a SAC agent for 100k transitions and use its replay buffer as the available offline dataset. Each observation consists of a centered, grayscale image of the pendulum, cropped to focus on the rod, resized to $64\times 64$, and stacked over the last four frames to provide temporal context. We train a residual convolutional autoencoder to obtain a compact latent representation of these stacked frames with latent dimension 32. In addition to the standard reconstruction loss, we introduce a latent-space dynamics loss by training an auxiliary dynamics predictor that maps the current latent and action to the next latent. This encourages the learned representation to better reflect the system’s underlying dynamics. After training the autoencoder, we apply our MPDiffuser framework in this latent space and compare its performance against Decision Diffuser, Planner-only, and Diffuser baselines. As summarized in Table 5, our approach achieves higher average returns, demonstrating that MPDiffuser can scale to visual domains and show superior performance even when operating on a learned latent space.
@@ 4.4 Robot Locomotion with Unitree Go2
To assess real-world applicability, we evaluate our framework on quadruped locomotion using the Unitree Go2. Experiments are done in IsaacLab (mittal2023orbit) using configurations from the official Unitree repository. The state includes base angular velocity, projected gravity, and joint angles and velocities. Default domain randomization parameters is applied inducing stochasticity in dynamics. The reward promotes accurate velocity tracking via an exponentially decaying penalty on tracking error, while the cost activates when the gravity projection’s $z$-coordinate exceeds 0.95-0.95, encouraging parallel torso with a safety budget of $10$. A PPO (schulman2017proximal) policy is trained for $1000$ epochs to track constant velocity commands. As our dataset, we use $5000$ rollouts from this policy at four training snapshots (epochs $100$, $400$, $700$, and $1000$).
Figure 10: Unitree Go2 quadruped robot walking.
In Table 6 we report average reward and cost computed over 1250 independent trials. MPDiffuser achieves the highest performance while remaining under the cost limit, highlighting the benefit of alternating planner–dynamics updates. In contrast, single-model baselines either generate unsafe trajectories or suffer from degraded returns due to their inability to balance task fidelity with dynamic feasibility. These results demonstrate that our compositional sampling strategy is crucial for reliable deployment in safety-critical locomotion tasks.
@@
Table 6: MPDiffuser matches performance, while maintaining safety. Performance of baseline methods and MPDiffuser on Unitree Go2 locomotion in simulation. We report normalized returns (relative to the dataset average) and normalized costs (relative to the cost budget).
Finally, we validate our approach on the Unitree Go2 quadruped, running fully onboard with a Jetson Orin. Due to limited compute, we use single-sample inference and system-level optimizations (see Section O). The robot tracks a constant velocity command of $[0.5,0,0]$, compared against a PPO policy trained for $1000$ epochs. Since direct velocity measurements are unavailable, a small MLP estimates velocity from joint states. As shown in Figure 10, PPO overshoots the target, while MPDiffuser closely matches it. The PPO policy achieves $0.74\,\text{m/s}$, whereas MPDiffuser maintains $0.55\,\text{m/s}$, near the commanded $0.5\,\text{m/s}$. Overall, MPDiffuser attains a normalized return of $1.02$ versus $0.98$ for PPO, both with zero cost. Consequently, this experiment demonstrates the practicality of MPDiffuser for real-world control problems.
@@ 5 Discussion
We introduced Model Predictive Diffuser (MPDiffuser), a model-based diffusion framework that composes planner, dynamics, and ranker modules to synthesize task-aligned and dynamically feasible behaviors from offline data. By interleaving planner and dynamics updates, our sampling scheme improves both fidelity to demonstrations and consistency with system dynamics, leading to state-of-the-art performance across unconstrained (D4RL) and constrained (DSRL) benchmarks, as well as real-world robotic deployment. While our focus has been on the offline setting, future work could explore extending MPDiffuser to online decision-making by leveraging the dynamics module for exploration or adaptive control. Another promising direction is scaling our framework to complex, high-dimensional sensory domains (e.g., vision-based control) by performing diffusion in compact latent spaces similar to xie2025latent, as preliminarily demonstrated in Sec. 4.3. Moreover, we aim to extend the framework across multiple environments and augment the dynamics model with cross-domain data, similar in spirit to world models (ha2018world).
Ethics Statement: This work does not raise any ethical concerns.
Reproducibility Statement: The benchmark datasets (D4RL, DSRL) are publicly available. For custom datasets, we describe the generation procedure in the respective sections. We also release a codebase that fully implements our method.
@@ Appendix A Related Work
Diffusion Model Based Control: Diffusion models have recently been applied to a wide range of decision-making and control problems. Early works such as pearce2023imitating, carvalho2023motion, and luo2025generative explored imitation learning and motion planning, showing that diffusion priors can generate smooth and diverse trajectories. In reinforcement learning, several approaches employ diffusion at the action level, where a single action is generated conditioned on the current state. For example, lu2023contrastive introduce Q-guided sampling and demonstrate strong reward performance. However, complex tasks with constraints and multiple objectives often require reasoning over longer horizons. To this end, trajectory-level diffusion has been adopted in offline RL settings (janner2022planning; ajay2023conditional), as well as for policy learning in robotics (chi2023diffusion; huang2025diffuse). These methods underscore the flexibility of diffusion-based formulations, as trajectory-level modeling captures long-term dependencies, composes behaviors observed in data, and accommodates constraints more effectively than single-step action generation.
@@
Dynamics-Aware Diffusion for Feasible Planning While diffusion models can effectively capture the distribution of state–action trajectories, generating trajectories that are dynamically feasible remains a fundamental challenge. Existing trajectory diffusion methods janner2022planning; ajay2023conditional synthesize rollouts directly in data space without enforcing the underlying system dynamics. As shown in follow-up studies zhou2025diffusion and corroborated by our results (Figure 2), this often yields trajectories that deviate from true transition structures—demonstrating that producing perfectly dynamically consistent sequences with diffusion models is inherently difficult. Several recent works attempt to alleviate this issue through inverse dynamics models (IDMs). For instance, ajay2023conditional first diffuse state sequences and infer actions via a learned IDM, but such trajectories are often unrealizable under true dynamics. Similarly, luo2025generative employ a related strategy for long-horizon planning and report frequent failure cases due to imperfect inverse dynamics. In contrast, our framework never relies on a single inverse-dynamics mapping: we explicitly separate planning from feasibility and correct the state evolution at every diffusion step using a dedicated dynamics diffusion model. This eliminates the brittle dependence on IDMs and keeps the trajectory close to the dynamics manifold throughout sampling. Safety-oriented extensions, such as umenberger, project states onto constraint manifolds during sampling but rely on the unrealistic assumption of a perfect inverse dynamics model for safety guarantees. MPDiffuser avoids such assumptions entirely: feasibility is enforced by a learned dynamics model operating at each diffusion timestep, yielding trajectories that satisfy constraints because the underlying state evolution is kept consistent with system transitions—not because an idealized inverse model is assumed. In the visual domain, xie2025latent apply an IDM over latent representations obtained from autoencoders; however, the resulting latent dynamics are often not well-posed, leading to severe degradation in control performance (Sec. 4.3). Our approach sidesteps this issue by maintaining a dedicated diffusion dynamics model even in latent space, ensuring that feasibility corrections remain well-defined and that visual rollouts do not drift into spurious latent transitions.
@@
Model Predictive Control and Diffusion-based Approximations: MPC is a leading optimization-based framework valued for its ability to optimize objectives under explicit constraints over finite horizons (rawlings2017model). Yet, solving its optimization online becomes intractable for complex models, intricate rewards, or nonconvex constraints. This has motivated approximate MPC, where offline solutions are used to train surrogates that approximate MPC behavior more efficiently (hertneck2018learning). Diffusion models have recently emerged as powerful generative surrogates. huang2024toward show that they can approximate MPC solutions with near-global optimality. However, diffusion models lack feasibility guarantees, creating a gap between generated trajectories and realized ones (zhao2024long). Several works aim to close this gap. zhou2025diffusion propose D-MPC with disjoint models for actions and dynamics, while our method integrates planning and dynamics correction within each diffusion step, simultaneously enforcing feasibility and improving trajectory fidelity. By integrating the dynamics model directly into each diffusion step, MPDiffuser incorporates dynamics feedback during sampling—whereas in D-MPC the dynamics model influences generation only indirectly through candidate ranking—so in the single-sample regime D-MPC’s dynamics model is effectively inert, while ours remains fully operational as an active component of the sampling process. romer2024diffusion adopt an alternating scheme, projecting trajectories onto feasible manifolds after each planner step via explicit optimization. Yet such projections are ill-posed within the diffusion process, since forward process breaks dynamic consistency. By contrast, our dynamics diffusion model learns the dynamics-induced manifold at every diffusion timestep, enabling feasibility enforcement in a distributionally consistent way during generation.
@@ Appendix B Hyperparameters and Model Architecture
Figure 11: Datasets for Deep Data-Driven Reinforcement Learning (D4RL) (d4rl).
@@
Figure 12: Datasets for Safe Reinforcement Learning (DSRL) (dsrl)
@@
  • Both the planner noise model $\epsilon_{\theta}^{\mathrm{pl}}$ and the dynamics noise model $\epsilon_{\theta}^{\mathrm{dyn}}$ are implemented as temporal U-Nets as proposed by janner2022planning. Each network consists of six repeated residual blocks, where each block contains two temporal convolutions, followed by group normalization wu2018group and a Swish activation swish. Conditioning inputs $y(\tau)$ and the initial state $x_{0}$ are first processed with a two-layer MLP and then injected into the U-Net through FiLM layers perez2018film.
  • We optimize $\epsilon_{\theta}$ and $f_{\phi}$ using Adam (kingma2014adam) with a learning rate of $2\times 10^{-4}$, a batch size of $64$, and $1\times 10^{6}$ training steps. We track an exponential moving average of the weights with decay $0.005$, which is employed for evaluation.

### v2

@@ 4.2 Constrained Offline Decision Making
Results on Standardized Benchmarks (DSRL): We evaluate our method on the DSRL benchmark (Liu et al., 2024), which includes safety-critical velocity and Safety Gym tasks. The objective is to maximize return while keeping cumulative cost below a specified budget. We compare against behavior cloning (BC-All), behavior cloning trained only on safe trajectories (BC-Safe), cost-regularized approaches COptiDICE (Lee et al., 2022), CPQ (Xu et al., 2022), BCQ-Lag (Xu et al., 2022), as well as transformer-based CDT (Liu et al., 2023). Our method, MPDiffuser, is tested using 16 samples with learned cost and reward functions parameterized as MLPs. For each task, all methods are evaluated under cost budgets of $20$, $40$, and $80$ reporting average normalized return and cost over $60$ trials per budget.
@@
Importance of Dynamic Feasibility for Ranker: We evaluate our approach on the classic Pendulum environment, modified with a hard velocity constraint requiring angular velocity to remain below $6.5\,\mathrm{m/s}$. We first train a standard soft actor-critic (SAC) (Haarnoja et al., 2018) agent for 100100k steps, which frequently violates the velocity constraint, and a safe SAC agent that penalizes constraint violations heavily. To construct the dataset, we use the replay buffer of the unsafe SAC agent and $300$ trajectories collected from the safe SAC agent. We then compare MPDiffuser with only planner based sampling and SafeDiffuser (Xiao et al., 2023) enforces safety by projecting sampled states within constraints during the diffusion process using a barrier-function.
@@
Table 5 shows success rates as a function of sampled trajectories, where success means stabilizing the pendulum upright without violating the velocity constraint. SafeDiffuser achieves substantially lower success rates than other methods, underscoring the necessity of dynamic feasibility for projection-based approaches to maintain safety. While planner initially improves with more samples, its performance later drops significantly because it often generates dynamically infeasible rollouts. With more samples, the chance of selecting such “hallucinated” trajectories that appear high-return but fail in practice increases. In contrast, MPDiffuser sustains high success rates as samples increase, highlighting robustness from our alternating sampling scheme.
Figure 7: Unitree Go2 quadruped robot walking.
@@ 4.3 Extending MPDiffuser to Visual Domains
To assess scalability to high-dimensional visual inputs, we conduct a preliminary proof-of-concept experiment on the Pendulum environment with image observations. We train a SAC agent for 100k transitions and use its replay buffer as the offline dataset. As observations we use centered grayscale images of the pendulum resized to $64\times 64$, and stacked over four frames for temporal context.
@@
We train a residual convolutional autoencoder to obtain a compact latent representation of these stacked frames with latent dimension 32. In addition to the standard reconstruction loss, we introduce a latent-space dynamics loss by training an auxiliary dynamics predictor that maps the current latent and action to the next latent. This encourages the learned representation to better reflect the system’s underlying dynamics. After training the autoencoder, we apply our MPDiffuser framework in this latent space and compare its performance against Decision Diffuser, Planner-only, and Diffuser baselines. As summarized in Table 6, our approach achieves higher average returns, demonstrating that MPDiffuser can scale to visual domains and show superior performance even when operating on a learned latent space.
@@ 4.4 Robot Locomotion with Unitree Go2
To assess real-world applicability, we evaluate our framework on quadruped locomotion using the Unitree Go2. Experiments are done in IsaacLab (Mittal et al., 2023) using configurations from the official Unitree repository. The state includes base angular velocity, projected gravity, and joint angles and velocities. Default domain randomization parameters is applied inducing stochasticity in dynamics. The reward promotes accurate velocity tracking via an exponentially decaying penalty on tracking error, while the cost activates when the gravity projection’s $z$-coordinate exceeds 0.95-0.95, encouraging parallel torso with a safety budget of $10$. A PPO (Schulman et al., 2017) policy is trained for $1000$ epochs to track constant velocity commands. As our dataset, we use $5000$ rollouts from this policy at four training snapshots (epochs $100$, $400$, $700$, and $1000$).
@@
Table 7: MPDiffuser matches performance while maintaining safety. Performance of baseline methods and MPDiffuser on Unitree Go2 simulation. Returns and costs are normalized relative to the dataset average and cost budget, respectively.
In Table 7 we report average reward and cost computed over $1,250$ trials. MPDiffuser achieves the highest performance while remaining under the cost limit, highlighting the benefit of alternating planner–dynamics updates. In contrast, single-model baselines either generate unsafe trajectories or suffer from degraded returns due to their inability to balance task fidelity with dynamic feasibility. These results demonstrate that our compositional sampling strategy is crucial for reliable deployment in safety-critical locomotion tasks.
Finally, we validate our approach on the Unitree Go2 quadruped, running fully onboard with a Jetson Orin. Due to limited compute, we use single-sample inference and system-level optimizations (see Sec. O). The robot tracks a constant velocity command of $[0.5,0,0]$, compared against a PPO policy trained for $1000$ epochs. Since direct velocity measurements are unavailable, a small MLP is trained to estimate velocity from joint states. As shown in Fig. 8, PPO overshoots the target, while MPDiffuser closely matches it. The PPO policy achieves $0.74\,\text{m/s}$, whereas MPDiffuser maintains $0.55\,\text{m/s}$, near the commanded $0.5\,\text{m/s}$. MPDiffuser attains a normalized return of $1.02$ versus $0.98$ for PPO, both with zero cost. Thus, this experiment confirms the practicality of MPDiffuser for real-world applications.
@@ 5 Discussion
We introduced Model Predictive Diffuser (MPDiffuser), a model-based diffusion framework that composes planner, dynamics, and ranker modules to synthesize task-aligned and dynamically feasible behaviors from offline data. By interleaving planner and dynamics updates, our sampling scheme improves both fidelity to demonstrations and consistency with system dynamics, leading to state-of-the-art performance across unconstrained (D4RL) and constrained (DSRL) benchmarks, as well as real-world robotic deployment. While our focus has been on the offline setting, future work could explore extending MPDiffuser to online decision-making by leveraging the dynamics module for exploration or adaptive control. Another promising direction is scaling our framework to complex, high-dimensional sensory domains (e.g., vision-based control) by performing diffusion in latent spaces similar to Xie et al. (2025), as preliminarily demonstrated in Sec. 4.3. Finally, we aim to extend the framework across multiple environments and augment the dynamics model with cross-domain data, similar in spirit to world models (Ha and Schmidhuber, 2018).
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Related Work
Diffusion Model Based Control: Diffusion models have recently been applied to a wide range of decision-making and control problems. Recent work such as Pearce et al. (2023); Carvalho et al. (2023), and Luo et al. (2025) explored imitation learning and motion planning, showing that diffusion priors can generate smooth and diverse trajectories. In reinforcement learning, several approaches employ diffusion at the action level, where a single action is generated conditioned on the current state. For example, Lu et al. (2023) introduce Q-guided sampling and demonstrate strong reward performance. However, complex tasks with constraints and multiple objectives often require reasoning over longer horizons. To this end, trajectory-level diffusion has been adopted in offline RL settings (Janner et al., 2022; Ajay et al., 2023; He et al., 2023), as well as for policy learning in robotics (Chi et al., 2023; Huang et al., 2025b). These methods underscore the flexibility of diffusion-based formulations, as trajectory-level modeling captures long-term dependencies, composes behaviors observed in data, and accommodates constraints more effectively than single-step action generation.
@@
Dynamics-Aware Diffusion for Feasible Planning While diffusion models can effectively capture the distribution of state–action trajectories, generating trajectories that are dynamically feasible remains a fundamental challenge. Existing trajectory diffusion methods (Janner et al., 2022; Ajay et al., 2023) synthesize rollouts directly in data space without enforcing the underlying system dynamics. As shown in follow-up studies (Zhou et al., 2025) and corroborated by our results (Figure 2), this often yields trajectories that deviate from true transition structures—demonstrating that producing perfectly dynamically consistent sequences with diffusion models is inherently difficult. Several recent works attempt to alleviate this issue through inverse dynamics models (IDMs). For instance, Ajay et al. (2023) first diffuse state sequences and infer actions via a learned IDM, but such trajectories are often unrealizable under true dynamics. Similarly, Luo et al. (2025) employ a related strategy for long-horizon planning and report frequent failure cases due to imperfect inverse dynamics. In contrast, our framework never relies on a single inverse-dynamics mapping: we explicitly separate planning from feasibility and correct the state evolution at every diffusion step using a dedicated dynamics diffusion model. This eliminates the brittle dependence on IDMs and keeps the trajectory close to the dynamics manifold throughout sampling. Safety-oriented extensions, such as Zhang et al. (2025), project states onto constraint manifolds during sampling but rely on the unrealistic assumption of a perfect inverse dynamics model for safety guarantees. MPDiffuser avoids such assumptions entirely: feasibility is enforced by a learned dynamics model operating at each diffusion timestep, yielding trajectories that satisfy constraints because the underlying state evolution is kept consistent with system transitions—not because an idealized inverse model is assumed. In the visual domain, Xie et al. (2025) apply an IDM over latent representations obtained from autoencoders; however, the resulting latent dynamics are often not well-posed, leading to severe degradation in control performance (Sec. 4.3). Our approach sidesteps this issue by maintaining a dedicated diffusion dynamics model even in latent space, ensuring that feasibility corrections remain well-defined and that visual rollouts do not drift into spurious latent transitions.
@@
Model Predictive Control and Diffusion-based Approximations: MPC is a leading optimization-based framework valued for its ability to optimize objectives under explicit constraints over finite horizons (Rawlings et al., 2017). Yet, solving its optimization online becomes intractable for complex models, intricate rewards, or nonconvex constraints. This has motivated approximate MPC, where offline solutions are used to train surrogates that approximate MPC behavior more efficiently (Hertneck et al., 2018). Diffusion models have recently emerged as powerful generative surrogates. Huang et al. (2025a) show that they can approximate MPC solutions with near-global optimality. However, diffusion models lack feasibility guarantees, creating a gap between generated trajectories and realized ones (Zhao et al., 2024). Several works aim to close this gap. Zhou et al. (2025) propose D-MPC with disjoint models for actions and dynamics, while our method integrates planning and dynamics correction within each diffusion step, simultaneously enforcing feasibility and improving trajectory fidelity. By integrating the dynamics model directly into each diffusion step, MPDiffuser incorporates dynamics feedback during sampling—whereas in D-MPC the dynamics model influences generation only indirectly through candidate ranking—so in the single-sample regime D-MPC’s dynamics model is effectively inert, while ours remains fully operational as an active component of the sampling process. Römer et al. (2025) adopt an alternating scheme, projecting trajectories onto feasible manifolds after each planner step via explicit optimization. Yet such projections are ill-posed within the diffusion process, since forward process breaks dynamic consistency. By contrast, our dynamics diffusion model learns the dynamics-induced manifold at every diffusion timestep, enabling feasibility enforcement in a distributionally consistent way during generation.
@@ Appendix B Hyperparameters and Model Architecture
Figure 9: Datasets for Deep Data-Driven Reinforcement Learning (D4RL) (Fu et al., 2021).
@@
Figure 10: Datasets for Safe Reinforcement Learning (DSRL) (Liu et al., 2024)
@@
  • Both the planner noise model $\epsilon_{\theta}^{\mathrm{pl}}$ and the dynamics noise model $\epsilon_{\theta}^{\mathrm{dyn}}$ are implemented as temporal U-Nets as proposed by Janner et al. (2022). Each network consists of six repeated residual blocks, where each block contains two temporal convolutions, followed by group normalization Wu and He (2018) and a Swish activation Ramachandran et al. (2017). Conditioning inputs $y(\tau)$ and the initial state $x_{0}$ are first processed with a two-layer MLP and then injected into the U-Net through FiLM layers Perez et al. (2018).
  • We optimize $\epsilon_{\theta}$ and $f_{\phi}$ using Adam (Kingma, 2014) with a learning rate of $2\times 10^{-4}$, a batch size of $64$, and $1\times 10^{6}$ training steps. We track an exponential moving average of the weights with decay $0.005$, which is employed for evaluation.